5. Microsoft currently pays a dividend of $2.30. The company’s goal is to increase dividends by 4% each year. You, the investor, require a rate of return of 12%.

5a. What is the current price you would pay now for a share in Microsoft?

5b. What will Microsoft’s dividend be in five years? What about six years?

5c. What will the price of a share in Microsoft be in five years?
Answer #1

5a). Current Price = [Current Dividend * (1 + g)] / [r - g]

= [$2.30 * (1 + 0.04)] / [0.12 - 0.04] = $2.392 / 0.08 = $29.90

5b). D5 = D0 * (1 + g)5 = $2.30 * (1 + 0.04)5 = $2.30 * 1.2167 = $2.80

D6 = D5 * (1 + g) = $2.80 * (1 + 0.04) = $2.91

5c). P5 = D6 / (r - g) = $2.91 / (0.12 - 0.04) = $2.91 / 0.08 = $36.38